OpenStudy (anonymous):

Suppose $9,400 is invested in an account at an annual interest rate of 3.2% compounded continuously. How long (to the nearest tenth of a year) will it take the investment to double in size?

OpenStudy (anonymous):

\[9400(1.032)^t=18800\] \[(1.032)^t=2\] \[t*\log{1.032}=\log{2}\] \[t=log{2}/log{1.032}\] t is about 22. (Little bigger than 22)

OpenStudy (anonymous):

Answer: 22 years.